OpenStudy (anonymous):

find the slope and the y-intercept of the line with the given equation. y=1/2x+6

OpenStudy (katrinakaif):

Slope = 1/2 Y intercept = 6

OpenStudy (anonymous):

how do you reach those answers ?

OpenStudy (anonymous):

i know m=slope , so 1/2 x and b= y intercept......just seems to easy to me

OpenStudy (katrinakaif):

Lol Well you got the reason for the answer with your own remark :P

OpenStudy (katrinakaif):

The equation is in y = mx + b Anything before the 'x' is the slope..or so called 'm' B is always the y - intercept